write an equation of the line that passes through the point and has the given slope.write the equation in slope intercept form
(3,0),m=-2

Answer by tw23279(14)

First, use the equation for point-slope form, which (y - y1) = m(x - x1), where (x1, y1) is a random point that would be on the line and m is the slope. And y and x are just variables.
Plug it in, (y - 0) = -2(x - 3)
Distribute.
And so you'll get y = -2x+6.